OpenStudy (anonymous):

What is the value of y in the solution to the following system of equations? 8x + 8y = -16 2x - 3y = -24

OpenStudy (anonymous):

use 8x+8y=-24 and make x the subject of the fomula, then x= -3 -y now substitute x=-3-y in -16 2x -3y=-24 and i think something is miss with -16 and 2x